Henry can paint the house in 4 hours, and Claude can paint the house in 3 hours. How long would it take the two of them to do the job together?
Answer by ewatrrr(24785)


Hi,
Let x represent the time it would take to do the job together
Question states. Per hour being the equalizer.
1/4 + 1/3 = 1/x Multiplying each side of the equation by 12x so as all denominators = 1
3x + 4x = 12
7x = 12
x = 1.714hr , time it would take to do the job together